OpenStudy (pounce0129):

if y=-8 when x=2, find x when y=32

OpenStudy (anonymous):

what why would you need to find x if you have it already

OpenStudy (anonymous):

trick question!!!!

OpenStudy (anonymous):

2/-8 = x/32 (cross multiply) (-8)(x) = (32)(2) -8x = 64 x = - 8 check... 2/-8 = -8/32 - 1/4 = - 1/4 (correct) ANSWER : x = - 8 when y = 32